OpenStudy (anonymous):

The height of a triangle is three inches less than the lenght of the base of the triangle. The area of the triangle is 90 in.2. Find the lenght of the triangle.

OpenStudy (anonymous):

H = B - 3 1st sentence Area of Triangle is 1/2 BH Substitute in for H from the first sentence. 90 = 1/2 B (B - 3) multiply both side by 2 180 = B(B - 3) 180 = B^2 - 3B make it = to 0 0 = B^2 - 3B - 180 Factor 0 = (B - 15)(B + 12) B = 15 or B = -12 You don't use the -12 because that can't be the Base length B = 15 H = 15 - 3 = 12